OpenStudy (anonymous):

Determine the value of eighteen r minus twenty-three when r equals 6. guys plz help

OpenStudy (anonymous):

ok, so its \[18r-23 (r=6)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, lets do it this way \[(18*6)-23\]

OpenStudy (anonymous):

\[18*6=108\]

OpenStudy (anonymous):

\[108-23=85\]

OpenStudy (anonymous):

the answer should be 85

OpenStudy (anonymous):

thanks man

OpenStudy (anonymous):

Anytime :)

OpenStudy (anonymous):

can u help me with this? Find the value of x to the second power plus x plus 1 when x equals 5.

OpenStudy (anonymous):

\[x ^{2}+x+1 (x=5)\]

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

plug everything in you get 5^2+5+1

OpenStudy (anonymous):

Simplify 30 + 1

OpenStudy (anonymous):

answer should be 31 :)

OpenStudy (anonymous):

thanks! :D

OpenStudy (anonymous):

np :) just pm me or tag me if you want my help again

OpenStudy (anonymous):

Last question To make a headband, Mrs. Jones needs 7 yards of ribbon. How many yards of ribbon does Mrs. Jones need to make twenty headbands? Let h represent the number of headbands. The algebraic expression for this is seven h.

OpenStudy (anonymous):

\[7h*20=140h\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so \{140h\} should be how many yards he needs

OpenStudy (anonymous):

\[140h\]

OpenStudy (anonymous):

Thanks so much man! :D

OpenStudy (anonymous):

No problem :) any time